Subject: APO 3.5m Users Committee minutes, 2/26/07

From: Suzanne Hawley

Submitted: Sat, 3 Mar 2007 22:51:07 -0800 (PST)

Message number: 1015 (previous: 1014, next: 1016 up: Index)

   Minutes, APO User's Committee meeting, 02/26/07

Attending: John Bally, Jon Fulbright, Jon Holtzman, Suzanne Hawley,
Michael Strauss, Remy Indebetouw, Mark Klaene, Scott Anderson

Absent: Al Harper, Bruce Gillespie, Russet McMillan

Minutes were taken by Michael Strauss and editted by Suzanne Hawley, since 
Bruce Gillespie was out of town.  Thanks Michael!

*******We started with representatives from each institution giving
feedback on their observing experiences over the last month:

   Princeton: Things are fine

   JHU: Things are fine

   Colorado: We're gearing up for the Pluto occultation observations
coming up.

   NMSU: There was some concern about the very long time it took
recently to change the wavelength setting on DIS, possibly
due to confusion because of an unfamiliar arc lamp?  Usually the
wavelength setting need not be done to high accuracy, so perhaps
just moving without checking could be an option in this situation.
It would be good to do some tests to determine the accuracy of the moves
so users could evaluate if not checking after the move in special
situations would be acceptable.

** ACTION: Mark Klaene will look into this situation and report back.

   Virginia: No news (completely weathered out this quarter!)

   UW: There were previous problems whereby ftp timed out early; this
has been fixed by setting the timeout period to an hour, thanks to
Fritz for dealing with this.

New business:

******* Suzanne asked whether it would be desireable to get SPICAM working
under TUI.  Doing so would take a few weeks of somebody's time.  The
concensus was that it would be a good thing to do, especially for new 
users (so they don't have to learn multiple interfaces).  However, 
SPICAM has a nice scripting language, and it wasn't clear whether 
TUI's scripting was as flexible (for example, it isn't clear if 
TUI's scripting allows do loops or the ability to include if statements).

Suzanne will follow up with Russell Owen about TUI scripting.  Users
are encouraged to provide feedback to their users reps about
integrating Spicam into TUI.


Mark Klaene discussed the telescope/instrument report (which was 
written by Bruce Gillespie). A few notes from the meeting are added.

***********0) Overview

This has been a relatively routine operational month at the site, 
with the exception of an unusual amount of snowfall.  The total 
snowfall for January recorded at Sunspot was the largest amount in 15 
years (35 inches), and the snow has continued falling into February. 
As is typical for this region, the storms have come in waves, wiping 
out observing for days at a time, followed by days of clear weather 
and decent observing conditions.  The site staff continue to give 
herculean effort to keep the site grounds, roads, and roof of the 3.5- 
m enclosure free of snow.  Otherwise, there have been fixable 
problems with DIS and SPIcam, and a Colorado staff member gave NIC- 
FPS an electronic tune-up.

   Despite the cold weather, there have been very few cold-weather
related problems at the mountain.

************1) Telescope

Other than routine maintenance and repairs, not much happened with 
the telescope during the period, and it performed very well.  There 
was some testing done related to CIF projects (see below), and a new 
version of TUI continued to be site-tested prior to release to the 

*************2) Instruments

DIS:  We experienced trouble with the red-side cryotiger, which was 
replaced.  There still seems to be intermittent trouble with the red- 
side cooling system--a clogged cold head is suspected, which will be 
serviced at the next opportunity.  A plan for dealing with the false 
signal caused by the red chip heater has been formulated but not yet 

The problems with the slitviewer shutter have not been seen this
month, after the original shutter was replaced.  However,
a new shutter with additional cold weather features has been
ordered, and should be even more robust.

Suzanne reported that tests of the new red chip confirm that the
fringing is much improved.  She will make these available to Jon
Holtzman for incorporation into the new DIS documentation.

SPIcam:  The ion pump was recently replaced, and also its controller.

NIC-FPS:  Stephane Beland (Colorado) made an on-site visit to work on 
various instrument electronic issues.  He was able to improve the 
false-signal gradient at the bottom of the chip, and determined a 
method to substantially reduce the read noise.

   We need to make sure that people are aware of these changes; John Bally
will send out a memo to apo35-general with the details.  The buffered
mode of reading reduces the readnoise from 40 e- to 30 e-.  The gain
has also changed, and Jon Holtzman will conduct some tests to 
quantify the new gain.

   John Bally is now the project scientist for NIC-FPS, taking over
from Jim Green, who has his hands full with COS.  John is planning a
visit to the site in a few weeks (as part of the Pluto occultation
campaign) where among other things he plans to implement a fast
postage stamp reading mode, to allow very fast readout for fast
variability observations.

John also needs to test the etalon piezos to confirm that they can
operate at cryogenic temperatures.  If this works well, the etalon can
simply be installed this upcoming summer; if not, there is a second
option of the etalon being installed in the dewar but at less than
cryogenic temperature via thermal isolation and a cold strap.

3) Engineering and CIF projects

The NA2 rotator upgrade project is proceeding, with a design review
planned at APO in early April.  The UVa team building TripleSpec will
be at the site also; delivery, integration, software and commissioning
issues will be discussed.  Triplespec will is currently planned to 
arrive at mountain in September, and may be available for shared-risk 
observing in the fourth quarter of this year.

Nick McDonald made measurements of the repeatability of the tertiary 
mirror rotation as part of the design studies for upgrading the 
tertiary mounting system.  He also has started assembling the rebuilt 
"old" telescope drive boxes that are intended as hot-swappable spares 
for the telescope.

The project to improve the calibration lamps on the telescope is 
moving along.  Several new test lamps have been ordered, and tests of 
a new flat-field white paint for the mirror covers yielded promising 

4) Miscellaneous

The anemometer at the top of the meteorological tower ceased to 
function, and the present weather conditions are not conducive to 
making repairs at this time--a broken fiber cable is suspected. 
Since wind is key weather factor for observing (or not observing), we 
are making do with hand-held anemometers while waiting for a new 
anemometer to be delivered and installed in a more convenient 
location as a permanent backup.  As of this writing, nearly all the 
parts for the backup anemometer have been received, and its 
installation will be accomplished shortly.

*************  AGILE update: Anjum Mukadam has a commissioning run in
a few weeks, if all goes well with the final instrument checkout at UW.

**************  User documentation
   Jon Holtzman has been working on this with Gabrelle Saurage. The new
DIS user's guide is now on the web; people have found it useful.  Jon
and Gabrelle are nearing a completed draft on NIC-FPS which will be
vetted by John Bally and then made available.  The plan is to do the
Echelle next as SPICAM has adequate documentation until it is
integrated into TUI.  Jon is eager for feedback from the user
community about this new documentation.  The website also needs 
some general cleaning up, to get rid of older, obsolete documents.

****************Summer shutdown plans

Mark Klaene gave a summary of the current plans for the summer shutdown:

   The plan is to recoat the primary at KPNO (the facility at Starfire has
some concerns about availability of environmental permits).  We've
asked for time in August if possible.  The entire shutdown will 
take roughly 3 weeks, the timing of which will depend on the 
availability of the KPNO facility.  While the primary is out, 
we'll service the optical support structure (under the mirror).
Other plans include:

   -Repaint the mirror covers.  There is a new flat-field paint which
should improve throughput in u by a factor of 2, but it is not very
durable, so we are still investigating options.

   -Recoat the tertiary.

   -Install the NA2 rotator upgrade (with TripleSpec in mind; it will
     require a larger torque load than the rotator can currently handle).

   We're also planning a 2-3 day shutdowns in May and June, to do two
     -Install new drive boxes
     -Carry out some work on the secondary actuators. 

   The next users committee meeting will be held Monday, March 19 at 
8:30 AM Pacific time.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1015 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to